OpenStudy (anonymous):

im stuck on this question x^2+12x+35 divided by 3x+15

OpenStudy (dumbcow):

Factor top and bottom....see what cancels

OpenStudy (anonymous):

Here.

OpenStudy (ash2326):

we have \[\frac{x^2+12x+35}{3x+15}\] let's take out 3 from the denominator we get \[\frac{x^2+12x+35}{3(x+5)}\] now let's divide \[x^2+12x+35\] by \[x+5\] ___________________________________ -5 | 1 12 35 -5 -35 1 7 0 so \[\frac{x^2+12x+35}{x+5}=x+7\] so we get \[ \frac{x+7}{3}\]

OpenStudy (preetha):

Briana, did you understand these explanations?